### v2

@@ 4 The Lower Tail
We first consider the case $\varepsilon\leq en^{-1/2}$. Set $k\coloneqq\hat{\varepsilon}^{-1}\ln(\hat{\varepsilon}^{2}n)/A$. Depending on whether ${\cal S}_{1}$ occurs or not and using the tower property of conditional expectation we obtain
@@
and the proof is completed in that case by noting that $A\leq\hat{\varepsilon}n$, see the computations around (26). In the rest of the proof we consider the case $\varepsilon>en^{-1/2}$, where $\hat{\varepsilon}=\varepsilon$. We set
@@
In order to complete the proof, we now consider the remaining case $A\leq k_{0}$ and $\varepsilon>en^{-1/2}$. First of all, note that if the dispersion process terminates in less than $k$ steps, then $j=X+1$ is such that $U_{\tau^{L}_{j}+t}=0$ for some $t\leq k$ and the process of unhappy particles does not go back above $\varepsilon n/8$ before it terminates. Recall also that we may assume $k<b$, so that $T_{n,M}\leq k$ and $X=i$ imply that $\tau_{i+1}^{L}\leq b$. If in addition we recall the definition of $\mathcal{S}_{i}$ in (27) and set
@@
$\mathbb{P}\big{(}T_{n,M}\leq k,$ $X<k_{0}/A\big{)}\leq\exp(-2^{-19}\varepsilon n)+\sum_{1\leq i<1+k_{0}/A}\mathbb{P}(T_{n,M}\leq k,X=i-1,{\cal S}_{i})$
@@
By applying Lemma 4.5 we obtain from $A\leq k_{0}=\exp(2^{-16}\varepsilon^{2}n)$ and $|\varepsilon|<1/9$
$\mathbb{P}\big{(}T_{n,M}\leq k,$ $X<k_{0}/A\big{)}\leq\exp(-2^{-19}\varepsilon n)+\frac{k_{0}}{A}\cdot\exp(-2^{-11}\varepsilon^{2}n)\leq\frac{2}{A}.$ (32)
Suppose now that $T_{n,M}\leq k$ and $X\geq k_{0}/A$. Recall that $\mathcal{M}_{i,j}$ is the event that the number of unhappy particles remains below $\varepsilon n/8$ for at least $j$ steps after $\tau_{i}^{L}$, and that $\tau_{i}^{L}\leq b$. Set
${\cal M}\coloneqq\bigcap_{1\leq i\leq k_{0}/A}{\cal M}_{i,\varepsilon^{-1}/2}.$
@@
($\mathbb{P}(\{T_{n,M}\leq b\}\cap{\cal M}^{c})\leq\frac{k_{0}}{A}\exp(2^{-11}\varepsilon^{2}n).$
@@
($\leq\mathbb{P}\left(\big{\{}T_{n,M}\leq k,X\geq k_{0}/A\big{\}}\cap\mathcal{M}\right)+\frac{k_{0}}{A}\exp(2^{-11}\varepsilon^{2}n)$
$=\frac{k_{0}}{A}\exp(2^{-11}\varepsilon^{2}n),$
where the last equality follows as the event $\mathcal{M}\cap\{X\geq k_{0}/A\}$ implies that the dispersion process runs for at least $\varepsilon^{-1}k_{0}/(2A)$ steps, and thus the probability of the events occurring simultaneously is zero. In addition, $k_{0}e^{-2^{-11}\varepsilon^{2}n}/A\leq A^{-1}$, whence we can conclude that
@@
Together with (32), this shows $\mathbb{P}(T_{n,M}\leq\varepsilon^{-1}k_{0}/(2A))\leq 3A^{-1}$ for $A\leq k_{0}$, completing the proof. ∎

### v3

@@ 4 The Lower Tail
We first consider the case $\varepsilon\leq en^{-1/2}$. Set $k\coloneqq\hat{\varepsilon}^{-1}\ln(\hat{\varepsilon}^{2}n)/A\geq 1$ (as otherwise there is nothing to show). Depending on whether ${\cal S}_{1}$ occurs or not and using the tower property of conditional expectation we obtain
@@
and the proof is completed in that case by noting that $A\leq\hat{\varepsilon}n$. In the rest of the proof we consider the case $\varepsilon>en^{-1/2}$, where $\hat{\varepsilon}=\varepsilon$. We set
@@
In order to complete the proof, we now consider the remaining case $A\leq k_{0}$ and $\varepsilon>en^{-1/2}$. Note that if the dispersion process terminates in at most $k$ steps, then $U_{\tau^{L}_{X+1}+t}=0$ for some $t\leq k$ and moreover, the process of unhappy particles does not go above $\varepsilon n/8$ after time $\tau_{X+1}^{L}$ before it terminates. Recall also that we may assume $k<b$, so that $T_{n,M}\leq k$ implies $\tau_{X+1}^{L}\leq b$. If in addition we recall the definition of $\mathcal{S}_{i}$ in (26) and set
@@
$\mathbb{P}\big{(}T_{n,M}\leq k,$ $~{}X<k_{0}/A\big{)}\leq\exp(-2^{-19}\varepsilon n)+\sum_{1\leq i<1+k_{0}/A}\mathbb{P}(T_{n,M}\leq k,X=i-1,{\cal S}_{i})$
@@
By applying Lemma 4.5 we obtain from $A\leq k_{0}=e^{2^{-16}\varepsilon^{2}n}$ and $|\varepsilon|<1/9$
$\mathbb{P}\big{(}T_{n,M}\leq k,$ $X<k_{0}/A\big{)}\leq\exp(-2^{-19}\varepsilon n)+\frac{2k_{0}}{A}\cdot\exp(-2^{-11}\varepsilon^{2}n)\leq\frac{3}{A}.$ (31)
Suppose now that $T_{n,M}\leq k$ and $X\geq k_{0}/A$. Let again $\mathcal{M}_{i,j}$ be the event that the number of unhappy particles remains below $\varepsilon n/8$ for at least $j$ steps after $\tau_{i}^{L}$, and that $\tau_{i}^{L}\leq b$. Set
${\cal M}\coloneqq\bigcap_{1\leq i<1+k_{0}/A}{\cal M}_{i,\varepsilon^{-1}/2}.$
@@
($\mathbb{P}(\{T_{n,M}\leq b\}\cap{\cal M}^{c})\leq\frac{2k_{0}}{A}\exp(-2^{-11}\varepsilon^{2}n).$
@@
($\leq\mathbb{P}\left(\big{\{}T_{n,M}\leq k,X\geq k_{0}/A\big{\}}\cap\mathcal{M}\right)+\frac{2k_{0}}{A}\exp(-2^{-11}\varepsilon^{2}n)$
$=\frac{2k_{0}}{A}\exp(-2^{-11}\varepsilon^{2}n),$
where the last equality follows as the event $\mathcal{M}\cap\{X\geq k_{0}/A\}$ implies that the dispersion process runs for more than $\varepsilon^{-1}k_{0}/(2A)$ steps, and thus the probability of the events occurring simultaneously is zero. In addition, $2k_{0}e^{-2^{-11}\varepsilon^{2}n}/A\leq 2/A$, whence we can conclude that
@@
Together with (31) this shows $\mathbb{P}(T_{n,M}\leq\varepsilon^{-1}k_{0}/(2A))\leq 5A^{-1}$ for $A\leq k_{0}$, completing the proof. ∎